# Compressing Chain-of-Thought in LLMs via Step Entropy

### v1

@@ Abstract
Large Language Models (LLMs) using Chain-of-Thought (CoT) prompting excel at complex reasoning but generate verbose thought processes with considerable redundancy, leading to increased inference costs and reduced efficiency. We introduce a novel CoT compression framework based on step entropy, a metric that quantifies the informational contribution of individual reasoning steps to identify redundancy. Through theoretical analysis and extensive empirical validation on mathematical reasoning benchmarks, we demonstrate that steps with low entropy are indeed highly redundant. Our experiments reveal that an astonishing 80% of low-entropy intermediate steps can be pruned with minor degradation in the final answer accuracy across DeepSeek-R1-7B, 14B and Qwen3-8B. This finding sharply contrasts with random or high-entropy pruning, which severely impairs reasoning performance. Building on this, we propose a novel two-stage training strategy combining Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) reinforcement learning. This approach enables LLMs to autonomously learn to generate compressed COTs during inference by strategically incorporating [SKIP] tokens. Our method significantly enhances LLM inference efficiency while rigorously preserving accuracy, offering profound implications for practical LLM deployment and a deeper understanding of reasoning structures. The code and data are released in https://github.com/staymylove/COT˙Compresstion˙via˙Step˙entropy.
Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex reasoning tasks, particularly when employing techniques like Chain-of-Thought (COT) (Wei et al. 2022). By generating explicit intermediate reasoning steps, often referred to as ”slow thinking”, Large Reasoning Model (LRM) such as the DeepSeek-R1 (Guo et al. 2025) Series and Qwen3 (Yang et al. 2025) significantly enhance performance on multi-step problems in domains like mathematics, coding and symbolic logic. This process allows the model to break down complex problems into more manageable components, leading to more reliable and accurate outcomes.
@@
However, a notable drawback of current slow thinking COT implementations is the considerable redundancy often present within the generated thought processes (Deng et al. 2023; Zhong et al. 2025a). These verbose reasoning paths, while thorough, lead to increased inference latency, higher computational costs, and diminished overall efficiency. As models become larger and are deployed at scale, these inefficiencies present a significant bottleneck for practical applications.
@@
To mitigate this, prior research has explored several compression strategies. One prominent direction focuses on making the CoT process implicit or latent, finetuning the model to internalize reasoning steps without verbalizing them (Deng, Choi, and Shieber 2024; Hao et al. 2024) or dynamically compressing them in latent space (Tan et al. 2025). Other work has focused on compressing the reasoning chain at different granularities, from pruning tokens in the input context (Li et al. 2023), enabling controllable token-level skipping during generation (Xia et al. 2025), to chunk-based compression (Wang et al. 2025b). While these methods improve efficiency, they do not offer a principled way to identify and remove entire reasoning steps that are semantically redundant.
@@
Intuitively, when humans tackle complex problems, they record only key milestones, omitting obvious thoughts. Recent work has sought to teach LLMs a similar ability to skip steps (Liu et al. 2024; Jiang, Li, and Ferraro 2025) or tune for length-compressible CoTs (Ma et al. 2025b). However, a fundamental question persists: how can we systematically identify which steps in a reasoning chain are crucial versus superfluous?
@@
Building on this validation, we introduce a two-stage training strategy combining Supervised Fine-Tuning (SFT) with Group Relative Policy Optimization (GRPO) (Shao et al. 2024) that enables models to autonomously generate compressed reasoning trajectories. The SFT stage teaches models to predict when to use [SKIP] tokens based on entropy-compressed training data, while GRPO optimizes a composite reward function balancing accuracy, compression ratio, and response length. Our trained models achieve 35-57% token reductions while maintaining or improving accuracy, demonstrating that LLMs can learn to perform efficient reasoning without sacrificing quality.
@@
This paper is structured as follows: Related Work reviews LLM reasoning with reinforcement learning and CoT compression techniques. Section 3 presents our entropy-based CoT compression methodology, including step entropy formulation, pruning strategy, and two-stage training for autonomous compression. Section 4 provides experimental validation across multiple benchmarks and models, establishing optimal pruning ratios and demonstrating our training approach’s effectiveness.
Figure 1: Comprehensive Performance of Chain-of-Thought Compression via Step Entropy. (a) Accuracy vs. Mask Ratio on 50 samples from DeepScaleR. This plot illustrates the impact of different pruning strategies (Random, High-Entropy Steps, Low-Entropy Steps) on final answer accuracy as the mask ratio of intermediate COT steps increases. Note that masking up to 80% of low-entropy steps maintains complete COT accuracy (74%), while high-entropy and random masking lead to significant performance degradation. (b) Accuracy vs. Tokens Usage Ratio on Other Benchmarks. This plot compares the accuracy and token usage ratio of the Full COT against our Compressed COT (80% low-entropy steps pruning) across Math500, AIME 2024, and AIME 2025 datasets of DeepSeek-R1-7B.
Related Work
LLM Reasoning with Reinforcement Learning
Reinforcement Learning (RL) has emerged as a powerful paradigm for enhancing the reasoning capabilities of Large Language Models (LLMs). Recent advancements, such as those demonstrated in (Shao et al. 2024) and (Xie et al. 2025), showcase RL’s efficacy in refining LLMs’ ability to tackle complex reasoning tasks. Furthermore, strategies involving ”long COT” (Yeo et al. 2025) and ”slow thinking” (Zhong et al. 2025b) (which involves extending inference time) (Comanici et al. 2025; Guo et al. 2025; OpenAI 2025) have been shown to significantly improve LLM reasoning performance by allowing for more elaborate and deliberate thought processes.
However, the increased length and computational overhead associated with these verbose COTs have led to concerns regarding efficiency. Research by (Wang et al. 2025a; Cuadron et al. 2025; Sui et al. 2025) highlights the phenomenon of ”overthinking,” where excessively long COTs can paradoxically lead to diminished efficiency without proportional gains in accuracy. This emphasizes the need for methods that can optimize the length and content of reasoning trajectories.
COT Compression and Latent Reasoning
@@
Explicit compression methods aim to shorten the generated text at various granularities. At the finest level, some works enable controllable token-level skipping (Xia et al. 2025) or explore the information-theoretic minimum number of tokens required for a solution (Lee, Che, and Peng 2025). At a coarser grain, R1-Compress introduces chunk-based compression and search (Wang et al. 2025b). Other strategies use length-constrained tuning, integrating penalties into RL reward functions (Shen et al. 2025; Hou et al. 2025) or developing specific architectures for length-compressible CoTs like CoT-Valve (Ma et al. 2025b). Our work advances this line by proposing a more semantically grounded approach: we operate at the step level, arguing it better mimics human cognition (skipping entire thoughts, not words). Furthermore, our method is guided by an explicit information-theoretic signal—step entropy—teaching the model not just to be shorter, but to selectively discard what is verifiably uninformative.
An alternative, more radical approach is to make reasoning implicit or latent. Methods like iCOT (Deng, Choi, and Shieber 2024) and COCONUT (Hao et al. 2024) fine-tune models to internalize reasoning steps, while others use knowledge distillation to embed the process in the model’s hidden states (Deng et al. 2023). More recently, dynamic latent compression performs reasoning entirely within these hidden states, avoiding explicit generation altogether (Tan et al. 2025). While these latent strategies offer maximum efficiency, they sacrifice the critical interpretability and verifiability of an explicit CoT. Our work carves a distinct path by focusing on optimizing the explicit reasoning chain, preserving its benefits while drastically improving its efficiency.
CoT Compression via Step Entropy
@@
$I(S_{j};A|\bar{S}_{j})$ $=I(S_{j};A|S_{1},...,S_{j-1},S_{j+1},...,S_{N})$ (3)
$=H(S_{j}|\bar{S}_{j})-H(S_{j}|\bar{S}_{j},A)$ (4)
$\leq H(S_{j}|\bar{S}_{j})=H(S_{j}|S_{<j},S_{>j})$ (5)
$=H(S_{j}|S_{<j})-I(S_{j};S_{>j}|S_{<j})$ (6)
Since $I(S_{j};S_{>j}|S_{<j})\geq 0$, we have $I(S_{j};A|\bar{S}_{j})\leq H(S_{j}|S_{<j})$. This result demonstrate that when the entropy of step $S_{j}$ is low, the conditional mutual information of $S_{j}$ and the final answer $A$, which implies that the relation of $S_{j}$ and $A$ is minor.
Now let’s assume that there are K steps with low entropy, $\tilde{S}=S_{k_{0}},S_{k_{1}},...,S_{K}$, then we have
Without loss of generality, we assume the indices $k_{i}$ are arranged in descending order, i.e., $k_{i}<k_{i-1}$. Therefore, we could split the sequence $C/[S_{k_{0}},...,S_{k_{i}}]$ into $S_{<k_{i}}$ and $S_{>k_{i}}/[S_{k_{0}},...,S_{k_{i-1}}]$. Now consider the item with $k_{i}$:
$=H(S_{k_{i}}|S_{<k_{i}},(S_{>k_{i}}/[S_{k_{0}},...,S_{k_{i-1}}]))$ (9)
$\leq H(S_{k_{i}}|S_{<k_{i}})$ (10)
Therefore, we conclude that
$\leq\sum_{i=0}^{K}H(S_{k_{i}}|S_{<k_{i}})$ (12)
The result denotes that steps $S_{k_{0}},S_{k_{1}},...,S_{K}$ could have minor relation to the final solution $A$. This observation implies that, a step with low entropy suggesting the deterministic thinking, which has minor relation to the finally solution, thus denotes such step could be less informative, and potentially redundant.
Figure 2: The case of Low-Entropy steps pruning strategy for COT compression, and replacing each selected low-entropy step with a special ‘[SKIP]’ token.
Low-Entropy Steps Pruning Strategy for COT Compression
Based on the theoretical foundation established above, we propose a practical CoT compression approach that selectively removes low-entropy steps while preserving the essential reasoning structure. Our method operates on the principle that steps with low entropy are more likely to be redundant and can be safely pruned without significantly impacting the final answer quality.
@@
    Entropy-Based Pruning: We rank all steps in ascending order of their entropy scores and identify the $\kappa\times N$ lowest-entropy steps for pruning, where $\kappa$ is the pruning ratio hyperparameter. The compressed CoT $C^{\prime}$ is constructed by replacing each selected low-entropy step with a special ‘[SKIP]’ token, while preserving high-entropy steps in their original form. The example of this process is shown in Figure 2.
@@
    Inference with Compressed CoT: The compressed sequence $C^{\prime}$ is concatenated with the user query and the </think> delimiter to prompt the model to generate only the final answer, as illustrated in Figure 3.
@@
This approach allows us to systematically compress CoT sequences while maintaining reasoning coherence. The pruning ratio $\kappa$ provides a flexible control mechanism for balancing compression efficiency and answer quality, with optimal values determined empirically across different datasets and problem types. The upper bound of steps pruning ratio $\kappa$ will be discussed in the experiment section.
Figure 3: Inference pipeline using compressed thinking COT. The compressed CoT sequence is concatenated with the user query in the prompt context. The </think> delimiter signals the model to generate only the final answer without additional reasoning.
Experiments
We present comprehensive experiments validating our entropy-based CoT compression method. We first establish the optimal pruning ratio through controlled experiments, then demonstrate the effectiveness and generalizability of our approach across multiple benchmarks and model sizes. Finally, we compare our step-level compression strategy against token-level alternatives to justify our methodological choices. Finally, we propose a two-stage training method to make LLM to learn to generate the compressed reasoning trajectory.
Method GSM8k Math500 AIME 2024 AIME 2025
ACC (%) Tokens ACC (%) Tokens ACC (%) Tokens ACC (%) Tokens
DeepSeek-R1-7B $80.36$ 390 545390\,545 $88.17$ 1 851 9171\,851\,917 $63.33$ 475 303475\,303 $35.71$ 546 097546\,097
DeepSeek-R1-7B (Our) $80.82$ 388180 ($\downarrow$0.6%) $88.17$ 1302114 ($\downarrow$29.7%) $56.67$ 302784 ($\downarrow$36.3%) $35.71$ 344135 ($\downarrow$37.0%)
DeepSeek-R1-14B $82.64$ 374 109374\,109 $84.37$ 1 426 8651\,426\,865 $65.52$ 462 445462\,445 $58.62$ 540 003540\,003
DeepSeek-R1-14B (Our) $84.00$ 367060 ($\downarrow$1.9%) $82.16$ 990486 ($\downarrow$30.6%) $58.62$ 261167 ($\downarrow$43.5%) $51.72$ 325262 ($\downarrow$39.8%)
Qwen3-8B $94.46$ 4 027 7934\,027\,793 $91.37$ 3 569 2473\,569\,247 $79.31$ 628 097628\,097 $76.92$ 597 073597\,073
Qwen3-8B (Our) $94.39$ 3373301 ($\downarrow$16.2%) $91.13$ 2604619 ($\downarrow$27.0%) $81.48$ 346007 ($\downarrow$44.9%) $76.00$ 351499 ($\downarrow$41.1%)
Table 1: Comparing the full COT baseline with our proposed step-entropy based pruning (Our) method, which prunes 80% of the lowest-entropy steps for DeepSeek-R1-7B, 14B and Qwen-8B. We conduct experiments to get the Pass@1 Accuracy(ACC)(%) and the number of Thinking Tokens (contains the Unicode characters) during the inference on GSM8k, Math500, AIME2024 and AIME2025.
Determining the Optimal Pruning Ratio
To identify the safe threshold for step pruning, we conduct a controlled experiment using 50 samples from DeepScaleR (Luo et al. 2025) dataset on DeepSeek-R1-7B. And we investigate the impact on final answer accuracy by pruning steps using three distinct strategies with a mask ratio varying from 10% to 100% (”no-thinking” mode (Ma et al. 2025a)), shown in Figure 1 (left).
  • Low-Entropy Steps Pruning: Steps with the lowest entropy scores are progressively removed.
  • High-Entropy Steps Pruning: Steps with the highest entropy scores are progressively removed.
  • Random Steps Pruning: Steps are removed at random, serving as a control.
The results, illustrated in Figure 1, strongly validate our hypothesis. With Low-Entropy Steps Pruning, we observe that final answer accuracy remains stable and unaffected even when up to 80% of the lowest-entropy steps are masked. Beyond this 80% threshold, accuracy begins to decline, eventually converging to the accuracy of the ”no-thinking” mode. This provides powerful evidence that a vast majority of low-entropy steps are indeed redundant. We found that with the best pruning strategy, we can prune up to 80% lowest-entropy steps (40% tokens redundancy) when not affecting the accuracy, shown in Figure 1 (right).
Conversely, with High-Entropy Steps Pruning, accuracy degrades immediately upon masking even a small fraction of steps. When the mask ratio exceeds 40%, performance drops below that of the ”no-thinking” mode, indicating that removing these critical, high-information steps is more detrimental than providing no reasoning at all. The Random Steps Pruning strategy’s performance falls between the two, beginning to decline at a 40% ratio.
Validating Low-Entropy Steps Pruning Strategy
Models and Datasets.
We use models from the DeepSeek-R1 series (7B and 14B) and Qwen3-8B (Yang et al. 2025), which are open-source Large Reasoning Models with strong mathematical reasoning capabilities. For generating the initial CoT trajectories and creating our training data, we use a combination of DeepScaleR (40k) (Luo et al. 2025) and OpenR1-Math (90k) (OpenR1-Math 2025). To test the effectiveness and generalizability of our compression method, we evaluate performance on several standard mathematical benchmarks: GSM8k (Cobbe et al. 2021), Math500 (Lightman et al. 2023), AIME2024 (dataset card AIME 2024) and AIME2025 (dataset card AIME 2025).
Performance and Generalizability on Benchmarks.
Table 1 presents comprehensive results comparing our compressed CoT method against full CoT baselines. Our approach consistently achieves substantial efficiency gains while maintaining or improving accuracy across all models. The DeepSeek-R1 series shows remarkable token reductions: 29.7-37.0% for the 7B model and 30.6-43.5% for the 14B model across mathematical benchmarks, with GSM8k showing slight accuracy improvements for both sizes. Notably, Qwen3-8B demonstrates the strongest performance with impressive token reductions of 16.2-44.9% while maintaining competitive accuracy and even achieving slight improvements on AIME 2024 (79.31%→81.48%). The cross-architecture consistency—spanning both DeepSeek-R1 and Qwen3 model families—demonstrates that step entropy is a robust and generalizable principle for identifying redundancy, independent of model architecture, size.
Scalability and Dataset Creation.
To ensure our method scales beyond controlled experiments, we further validated that the 80% low-entropy pruning strategy holds at a much larger scale. To validate this strategy at scale, we applied this compression pipeline to the entire DeepScaler (40k) and OpenR1-Math datasets (90k). The results in Table 2 confirm that even across these tens of thousands of examples, the accuracy of the statically compressed CoTs remains almost identical to that of the full CoTs. This large-scale validation not only proves the robustness of our method but also serves as the direct procedure for CoT Compression for Dataset Creation. The resulting dataset, consisting of (problem, compressed CoT) pairs, serves as the foundation for our training strategy.
Model DeepScaleR OpenR1-Math
Full CoT 63.74 48.39
Compressed COT 62.17 47.31
Figure 4: Comparing the accuracy of Our Method (via step-entropy) and Directly Masking Tokens (via token-entropy) across various thinking token mask ratios, with Full COT serving as the baseline.
Discussion of Our Method v.s. Directly Masking Tokens
A crucial aspect of our methodology is the decision to prune entire reasoning steps rather than individual tokens. To justify this, we compared our step-based pruning approach against a token-based pruning baseline, where we remove the lowest-entropy tokens from the thinking trace irrespective of the steps they belong to. The results, shown in Figure 4, are unequivocal. While our step-pruning method maintains baseline accuracy even after removing up to 40% of the total thinking tokens, the token-pruning approach leads to a sharp and immediate decline in performance. Accuracy drops significantly after just a 20% token mask ratio. This demonstrates that a reasoning step is the correct semantic unit for compression. Removing individual low-entropy tokens (e.g., common words or operators) can break the syntactic and semantic integrity of a critical reasoning step, rendering it incomprehensible to the model. In contrast, removing an entire low-entropy step preserves the structure of the remaining, more important steps, leading to a much more robust compression strategy.
Two-Stage Training Strategy

### v2

# Making Slow Thinking Faster: Compressing LLM Chain-of-Thought via Step Entropy

@@ Abstract
Large Language Models (LLMs) using Chain-of-Thought (CoT) prompting excel at complex reasoning but generate verbose thought processes with considerable redundancy, leading to increased inference costs and reduced efficiency. We introduce a novel CoT compression framework based on step entropy, a metric that quantifies the informational contribution of individual reasoning steps to identify redundancy. Through theoretical analysis and extensive empirical validation on mathematical reasoning benchmarks, we demonstrate that steps with low entropy are indeed highly redundant. Our experiments reveal that an astonishing 80% of low-entropy intermediate steps can be pruned with minor degradation in the final answer accuracy across DeepSeek-R1-7B, 14B and Qwen3-8B. This finding sharply contrasts with random or high-entropy pruning, which severely impairs reasoning performance. Building on this, we propose a novel two-stage training strategy combining Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) reinforcement learning. This approach enables LLMs to autonomously learn to generate compressed COTs during inference by strategically incorporating [SKIP] tokens. Our method significantly improves LLM inference efficiency while preserving accuracy, paving the way for more scalable LLM deployments and a better understanding of their internal reasoning. The code and data are released in https://github.com/staymylove/COT_Compresstion_via_Step_entropy.
1 Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex reasoning tasks, particularly when employing techniques like Chain-of-Thought (COT) (Wei et al., 2022). By generating explicit intermediate reasoning steps, often referred to as “slow thinking”, Large Reasoning Model (LRM) such as the DeepSeek-R1 (Guo et al., 2025) Series and Qwen3 (Yang et al., 2025) significantly enhance performance on multi-step problems in domains like mathematics, coding and symbolic logic. This process allows the model to break down complex problems into more manageable components, leading to more reliable and accurate outcomes.
@@
However, a notable drawback of current slow thinking COT implementations is the considerable redundancy often present within the generated thought processes (Deng et al., 2023; Zhong et al., 2025a). These verbose reasoning paths, while thorough, lead to increased inference latency, higher computational costs, and diminished overall efficiency. As models become larger and are deployed at scale, these inefficiencies present a significant bottleneck for practical applications.
@@
To mitigate this, prior research has explored several compression strategies. One prominent direction focuses on making the CoT process implicit or latent, finetuning the model to internalize reasoning steps without verbalizing them (Deng et al., 2024; Hao et al., 2024) or dynamically compressing them in latent space (Tan et al., 2025). Other work has focused on compressing the reasoning chain at different granularities, from pruning tokens in the input context (Li et al., 2023), enabling controllable token-level skipping during generation (Xia et al., 2025), to chunk-based compression (Wang et al., 2025a). While these methods improve efficiency, they do not offer a principled way to identify and remove entire reasoning steps that are semantically redundant.
@@
Intuitively, when humans tackle complex problems, they record only key milestones, omitting obvious thoughts. Recent work has sought to teach LLMs a similar ability to skip steps (Liu et al., 2024; Jiang et al., 2025) or tune for length-compressible CoTs (Ma et al., 2025b). However, a fundamental question persists: how can we systematically identify which steps in a reasoning chain are crucial versus superfluous?
@@
Building on this validation, we introduce a two-stage training strategy combining Supervised Fine-Tuning (SFT) with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) that enables models to autonomously generate compressed reasoning trajectories. The SFT stage teaches models to predict when to use [SKIP] tokens based on entropy-compressed training data, while GRPO optimizes a composite reward function balancing accuracy, compression ratio, and response length. Our trained models achieve 35-57% token reductions while maintaining or improving accuracy, demonstrating that LLMs can learn to perform efficient reasoning without sacrificing quality.
@@
2 Related Work
2.1 LLM Reasoning with Reinforcement Learning
Reinforcement Learning (RL) has emerged as a powerful paradigm for enhancing the reasoning capabilities of Large Language Models (LLMs). Recent advancements, such as those demonstrated in Shao et al. (2024) and Xie et al. (2025), showcase RL’s efficacy in refining LLMs’ ability to tackle complex reasoning tasks. Furthermore, strategies involving “long COT” Yeo et al. (2025) and “slow thinking” Zhong et al. (2025b) (which involves extending inference time) Comanici et al. (2025); Guo et al. (2025); OpenAI (2025) have been shown to significantly improve LLM reasoning performance by allowing for more elaborate and deliberate thought processes.
However, the increased length and computational overhead associated with these verbose COTs have led to concerns regarding efficiency. Wang et al. (2025b); Cuadron et al. (2025); Sui et al. (2025) highlight the phenomenon of “overthinking,” where excessively long COTs can paradoxically lead to diminished efficiency without proportional gains in accuracy. This emphasizes the need for methods that can optimize the length and content of reasoning trajectories.
2.2 COT Compression and Latent Reasoning
@@
Explicit compression methods aim to shorten the generated text at various granularities. At the finest level, some works enable controllable token-level skipping (Xia et al., 2025) or explore the information-theoretic minimum number of tokens required for a solution (Lee et al., 2025). At a coarser grain, R1-Compress introduces chunk-based compression and search (Wang et al., 2025a). Other strategies use length-constrained tuning, integrating penalties into RL reward functions (Shen et al., 2025; Hou et al., 2025) or developing specific architectures for length-compressible CoTs like CoT-Valve (Ma et al., 2025b). Our work advances this line by proposing a more semantically grounded approach: we operate at the step level, arguing it better mimics human cognition (skipping entire thoughts, not words). Furthermore, our method is guided by an explicit information-theoretic signal—step entropy—teaching the model not just to be shorter, but to selectively discard what is verifiably uninformative.
An alternative, more radical approach is to make reasoning implicit or latent. Methods like iCOT (Deng et al., 2024) and COCONUT (Hao et al., 2024) fine-tune models to internalize reasoning steps, while others use knowledge distillation to embed the process in the model’s hidden states (Deng et al., 2023). More recently, dynamic latent compression performs reasoning entirely within these hidden states, avoiding explicit generation altogether (Tan et al., 2025). While these latent strategies offer maximum efficiency, they sacrifice the critical interpretability and verifiability of an explicit CoT. Our work carves a distinct path by focusing on optimizing the explicit reasoning chain, preserving its benefits while drastically improving its efficiency.
Figure 1: Comprehensive Performance of COT Compression via Step Entropy. (a) Accuracy vs. Mask Ratio on 50 samples from DeepScaleR. This plot illustrates the impact of different pruning strategies (Random, High-Entropy Steps, Low-Entropy Steps) on final answer accuracy as the mask ratio of intermediate COT steps increases. Note that pruning up to 80% low-entropy steps maintains Complete COT accuracy (b) Accuracy vs. Tokens Usage Ratio on other benchmarks. This plot compares the accuracy and token usage ratio of the Full COT against our Compressed COT (80% low-entropy steps pruning) across Math500, AIME 2024, and AIME 2025 on DeepSeek-R1-7B.
3 CoT Compression via Step Entropy
@@ 3.1 Step Entropy
Lemma 1 (Entropy-Bounded Information).
Given a reasoning process where a sequence of steps $C=(S_{1},S_{2},\dots,S_{N})$ leads to a final answer $A$, the conditional mutual information $I(S_{j};A|\bar{S}_{j})$ between the step and the answer, conditioned on all other steps $\bar{S}_{j}=C\setminus\{S_{j}\}$, is bounded by the step entropy $H(S_{j}|S_{<j})$ of the step itself. Specifically:
$I(S_{j};A|\bar{S}_{j})\leq H(S_{j}|S_{<j})$
The formal proof is provided in Appendix B.1. This result demonstrate that when the entropy of step $S_{j}$ is low, the conditional mutual information of $S_{j}$ and the final answer $A$, i.e. the relation of $S_{j}$ and $A$ is minor.
Theorem 1 (Entropy-Bounded Information on Subset).
For any subset of $K+1$ steps $\tilde{S}=\{S_{k_{0}},S_{k_{1}},\dots,S_{k_{K}}\}\subseteq C$, the conditional information contribution of this subset to the final answer is bounded by the sum of the step entropies of the steps within it:
$I(\tilde{S};A|C\setminus\tilde{S})\leq\sum_{i=0}^{K}H(S_{k_{i}}|S_{<k_{i}})$
The formal proof is provided in Appendix B.2 The result denotes that steps $S_{k_{0}},S_{k_{1}},...,S_{k_{K}}$ could have minor relation to the final solution $A$. This observation implies that, a step with low entropy suggesting the deterministic thinking, which has minor relation to the finally solution, thus denotes such step could be less informative, and potentially redundant.
While step entropy theoretically serves as an upper bound on the mutual information between the CoT steps and the response, in practice, it suffers from bias introduced by step length. To mitigate this issue, we propose to use the length-normalized step entropy, as defined in Definition 1.
Definition 1 (Length-normalized Step Entropy).
Given a reasoning step $S_{i}$ consisting of $M_{i}$ tokens and the preceding context $S_{<i}$, we define the length-normalized step entropy as:
$H(S_{i}\mid S_{<i})=\frac{1}{M_{i}}\sum_{j=1}^{M_{i}}H(t_{i,j}\mid c_{i,j})$ (3)
where $t_{i,j}$ represents the $j$-th token of step $i$ and $c_{i,j}$ represents the context up to that token.
(a) The case of Low-Entropy steps pruning strategy for COT compression, and replacing each selected low-entropy step with a special ‘[SKIP]’ token.
(b) The compressed CoT sequence (<think></think>) is concatenated with the user query in the prompt context.
3.2 Low-Entropy Steps Pruning Strategy for COT Compression
Based on the observation established above, we propose a practical CoT compression approach that selectively removes low-entropy steps while preserving the essential reasoning structure. Our method operates on the principle that steps with low entropy are more likely to be redundant and can be safely pruned without significantly impacting the final answer quality.
@@ 3.2 Low-Entropy Steps Pruning Strategy for COT Compression
    Entropy-Based Pruning: We rank all steps in ascending order of their entropy scores and identify the $\kappa\times N$ lowest-entropy steps for pruning, where $\kappa$ is the pruning ratio hyperparameter. The compressed CoT $C^{\prime}$ is constructed by replacing each selected low-entropy step with a special ‘[SKIP]’ token, while preserving high-entropy steps in their original form. The example of this process is shown in Figure 2(a).
@@
    Inference with Compressed CoT: The compressed sequence $C^{\prime}$ is concatenated with the user query and the </think> delimiter to prompt the model to generate only the final answer, as illustrated in Figure 2(b).
@@
This approach allows us to systematically compress CoT sequences while maintaining reasoning coherence. The pruning ratio $\kappa$ provides a flexible control mechanism for balancing compression efficiency and answer quality, with optimal values determined empirically across different datasets and problem types. The upper bound of steps pruning ratio $\kappa$ will be discussed in Experiment 4.1.
3.3 Autonomous Compression via Two-Stage Training